# MasonTigers at SemEval-2024 Task 9: Solving Puzzles with an Ensemble of Chain-of-Thoughts

### v1

@@ Abstract
Our paper presents team MasonTigers submission to the SemEval-2024 Task 9 - which provides a dataset of puzzles for testing natural language understanding. We employ large language models (LLMs) to solve this task through several prompting techniques. Zero-shot and few-shot prompting generate reasonably good results when tested with proprietary LLMs, compared to the open-source models. We obtain further improved results with chain-of-thought prompting, an iterative prompting method that breaks down the reasoning process step-by-step. We obtain our best results by utilizing an ensemble of chain-of-thought prompts, placing $2^{nd}$ in the word puzzle subtask and $13^{th}$ in the sentence puzzle subtask. The strong performance of prompted LLMs demonstrates their capability for complex reasoning when provided with a decomposition of the thought process. Our work sheds light on how step-wise explanatory prompts can unlock more of the knowledge encoded in the parameters of large models.
MasonTigers at SemEval-2024 Task 9: Solving Puzzles with an Ensemble of Chain-of-Thoughts
@@ 1 Introduction
In recent years, Large Language Models have achieved impressive performance on several question-answering and language-understanding tasks when provided with appropriate prompting and conditioning Brown et al. (2020). However, complex reasoning abilities often present a challenge for these models. Task 9 of SemEval-2024 introduces a novel dataset called BrainTeaser Jiang et al. (2023) focused precisely on evaluating these reasoning capabilities with a set of complex puzzles and/or brainteasers. Such tasks involve solving word and sentence puzzles, which require multi-step inference and deduction. The dataset covers a diverse range of puzzle types including sequences, analogies, classifications, mathematical reasoning, inferences about implicit relationships, and more. Solutions frequently demand a chained application of knowledge and logic across multiple steps to uncover insights or concepts not directly stated in the problem description.
@@
Solving these elaborative reasoning problems remains an open challenge for natural language processing systems. We explore whether and how large language models can succeed on this task through prompting methodology. We employ proprietary models such as GPT-4 OpenAI (2023) and Claude 2.1 Anthropic (2023) through API - which have shown promising few-shot reasoning ability. We also use Mixtral Jiang et al. (2024), an open-source LLM that shows state-of-the-results in several language reasoning tasks. The prompting paradigm involves providing models with natural language descriptions that encode the reasoning process step-by-step Liu et al. (2021). We implement various prompting approaches for mapping puzzles to conditional text and systematically transforming reasoning into explanation chains. Our core method, chain-of-thought prompting Wei et al. (2022), iteratively breaks down the deduction into simplified logical steps.
@@
Experiments reveal that while zero-shot performance lags due to a lack of grounding, multi-step prompts can unlock substantial reasoning ability in models. Performance improves with more steps and more specificity in the prompt. While introducing few-shot prompting generates good results - the models do significantly better with chain-of-thought prompting. We experiment with several chains of thought and achieve mostly similar results with each attempt. To make a more empirically confident guess towards solving the puzzles we adopt an ensemble of these chains based on majority voting. Our approach achieves competitive performance, ranking 2nd on the word puzzle subtask and 13th on sentence puzzles.
2 Related Works
@@
Large Language Models have been used for more and more advanced and complicated tasks recently Raihan et al. (2023a, b); Goswami et al. (2023). They have also shown good reasoning abilities in several tasks. The task of solving puzzles and the BrainTeaser dataset Jiang et al. (2023) are respectively a novel task and a novel dataset in the NLP domain. Similar to their Multiple Choice Questions (MCQs) approach - a few datasets like MathQA Austin et al. (2021) have been compiled. However, these are intended for specific tasks and do not necessarily require deep reasoning to solve them; rather, sufficient domain knowledge is good enough. A similar work is done by Saeedi et al. (2020) where they investigate a task that combines natural language understanding and commonsense reasoning. They present deep learning architectures for distinguishing between sensible and nonsensical statements, achieving competitive results.
@@
In the field of reasoning task, chain of thought Wei et al. (2022) implies a logical sequence of connected ideas, fostering coherence and depth in responses. On the other hand, a tree of thought suggests branching out into various related ideas, offering a more comprehensive exploration of a topic. While few-shot prompting is effective for some tasks by providing examples to guide the model, it may have limitations in capturing the complexity of nuanced conversations. The optimal choice may involve a hybrid approach, where a few-shot prompt sets the initial Yao et al. (2023) context, and the model subsequently follows a chain or tree of thought to generate more contextually rich and coherent responses useful for reasoning tasks.
Tan (2023) shows the performance of LLM’s on the reasoning of Arithmetic word problems. It states that higher degrees of realization are associated with better overall accuracy on arithmetic problems. And chain of thought is really helpful in this aspect as it covers a variety of prompts to strengthen the reasoning.
Similarly, the paper Mo and Xin (2023) presents a new reasoning framework for large language models by addressing a gap in prior tree-based reasoning methods which overlooked inherent uncertainties in intermediate decision points made by models. Overall, the key innovation is leveraging uncertainty estimation locally within the models during tree reasoning to enable more precise problem-solving and reasoning.
3 BrainTeaser
The BrainTeaser dataset Jiang et al. (2023) is a question-answering benchmark designed to evaluate models’ ability for lateral thinking, i.e., to defy default commonsense associations and reason creatively. The dataset contains 1,100 multiple-choice questions divided into two subtasks - 627 sentence-based puzzles relying on narrative context and common phrases and 492 puzzles focused on the literal form and letters of words.
For a fair comparison with human performance, the dataset also provides a separate human evaluation set with 102 randomly sampled questions. Each question in BrainTeaser has one correct answer and three distractor choices, including the option "None of the above" to simulate creative reasoning failure.
Property Sentence Word
# Prompts 627 492
Avg. tokens 34.88 10.65
% of $>$30 tokens 48.32% 2.23%
Avg. answer length (tokens) 9.11 3.0
std. dev. [Choices] 2.36 0.52
Table 1: Key statistics of the BrainTeaser dataset partitions.
To prevent memorization of training data, the dataset also contains semantically and contextually reconstructed variants for every question while preserving the original reasoning process and answers. Statistics for the key properties of the dataset partitions are shown in Table 1.
During SemEval Task 9, during the dev phase, a total of 240 prompts (120 for both sentence and word puzzles) are provided, and during the test phase, a total of 216 prompts (120 for sentence and 96 for word puzzles) are provided.
@@ 4 Experiments
In our experiments, we focus on several prompting strategies by employing three state of the art models including proprietary models like GPT-4 OpenAI (2023) and Claude 2.1 Anthropic (2023) (accessed via API key) and one open-source model - Mixtral Jiang et al. (2024).
@@ 4.3 Chain-of-Thought
Ensemble of Chain-of-Thoughts
@@
This ensemble methodology with Chain of Thought prompting helps improve robustness to outlier examples and noise compared to using a single model. By prompting the ensemble components on different random question subsets, diversity is promoted to capture a more holistic representation of the overall data distribution. The voting also helps cancel out issues with single models latching onto spurious patterns. Overall, the ensemble approach with multiple Chain of Thought prompt sets provides a robust assessment strategy suited for the open-ended nature and diversity of the BrainTeaser puzzles.
@@ 5 Results
We analyze the performance of the three models - including GPT4 Turbo, Claude 2.1, and Mixtral. These models are tested with different types of prompts - regular and Chain-of-Thought (CoT), and with a varying number of examples, known as shots, ranging from zero to eight. Additionally, an ensemble method is applied to the eight-shot CoT prompting to see if it can further improve the models’ performance.
The results, shown in Table 2, reveal how the models performed under each condition. A human baseline with scores of 0.91 for both Sentence and Word puzzles in the test set is provided by the task organizers for comparison purposes.
@@
Table 2: Comparing the results generated by the models with different prompting strategies. [CoT] - denotes Chain-of-Thought. [E] - denotes Ensemble (as described in 4.3). [*] - denotes submission during the test phase on the LeaderBoard.
@@
GPT4 Turbo shows the best performance, especially with CoT prompting and an increasing number of shots. The model performs best with the eight-shot CoT prompting combined with the ensemble method ([E]), reaching the highest Sentence and Word scores of 0.93 and 0.95 in the test set, respectively. This shows that CoT prompting and the ensemble method significantly improve the model’s understanding and output.
Claude 2.1 also improves with CoT prompting and more shots. Its best scores were with the eight-shot CoT with the ensemble, achieving Sentence and Word scores of 0.86 and 0.95 in the test set, respectively. The asterisk (*) mark in Table 2 denotes our submission during the test phase.
Mixtral consistently gets better with more shots and CoT prompting. Its best results come with the eight-shot CoT and the ensemble technique, with Sentence and Word scores of 0.88 and 0.82 in the test set, respectively. Though it makes progress, Mixtral’s performance indicates inferiority compared to the other two models in this task.
@@
The results highlight the effectiveness of CoT prompting in boosting the performance of LLMa. This approach, especially when combined with more examples and the ensemble method, greatly improves models’ abilities to process and generate more accurate responses. GPT4 Turbo’s top performance is likely due to its advanced design, which makes the most of these strategies. On the other hand, Claude 2.1’s results point to the importance of model-specific adjustments.
6 Conclusion
@@
In this paper, we presented our team MasonTigers’ approach to SemEval-2024 Task 9 on solving puzzles using large language models. We explored various prompting strategies to guide the models, including zero-shot, few-shot, and chain-of-thought prompting. Our key method involved iteratively breaking down reasoning into simplified logical steps to decompose the complex deduction process.
@@
Experiments revealed promising results - while zero-shot performance was limited, providing explanatory prompts substantially improved the models’ reasoning abilities. Performance increased with more prompt specificity and steps. Our best results came from an ensemble approach applying majority voting across multiple chain-of-thought prompts.
@@
Ultimately, our system achieved highly competitive rankings on the leaderboard, placing 2nd on word puzzles and 13th on sentence puzzles. The strong capability unlocked through guided prompting highlights these models’ latent reasoning potential when given a structured thought process. Our work sheds light on how explanatory chains can elicit more of the knowledge within large language model parameters.
Limitations
@@
A few key limitations remain to be addressed in future work. First, constructing effective prompts requires extensive human effort and insight - automating this prompting process could improve scalability. Additionally, performance still lags behind human levels, suggesting room for advancement. Architectural constraints related to long-term memory and reasoning likely need to be overcome. Finally, our approach focused narrowly on the given puzzles rather than teaching broader inferential skills - developing more generalizable reasoning through prompts is an open challenge. Overall, this task highlights the promises and limitations of large language models in advanced reasoning.

### v2

# MasonTigers at SemEval-2024 Task 9: Solving Puzzles with an Ensemble of Chain-of-Thought Prompts

@@ Abstract
This paper presents the MasonTigers’ submission to the SemEval-2024 Task 9 which provides a dataset of puzzles for testing natural language understanding. We employ large language models (LLMs) to solve this task through several prompting techniques. We show that zero-shot and few-shot prompting with proprietary LLMs outperform open-source models. Results are further improved with chain-of-thought prompting. We obtain our best results by utilizing an ensemble of chain-of thought prompts, ranking $2^{nd}$ in the word puzzle sub-task and $13^{th}$ in the sentence puzzle sub-task.
MasonTigers at SemEval-2024 Task 9: Solving Puzzles with an Ensemble of Chain-of-Thought Prompts
@@ 1 Introduction
In recent years, LLMs have achieved impressive performance on several question answering and language understanding tasks when provided with appropriate prompting Brown et al. (2020). However, complex reasoning abilities often present a challenge for these models. SemEval-2024 Task 9 Jiang et al. (2024b) introduces a novel dataset called BrainTeaser Jiang et al. (2023) which includes a set of complex puzzles and brainteasers. Such tasks involve solving word and sentence puzzles, which require multi-step inference and deduction. The dataset covers a diverse range of puzzle types including sequences, analogies, classification, mathematical reasoning, inferences about implicit relationships, and more. Solutions frequently demand a chained application of knowledge and logic across multiple steps to uncover insights or concepts not directly stated in the problem description.
@@
Solving these elaborate reasoning problems is a challenging scenario for NLP systems. We explore whether and how LLMs can succeed on this task. We employ proprietary models such as GPT-4 OpenAI (2023) and Claude 2.1 Anthropic (2023) through APIs. These models have shown promising few-shot reasoning ability. We also use Mixtral Jiang et al. (2024a), an open-source LLM that shows state-of-the-results in several language reasoning tasks. The prompting paradigm involves providing models with natural language descriptions that encode the reasoning process step-by-step Liu et al. (2021). We implement various prompting approaches for mapping puzzles to conditional text and systematically transforming reasoning into explanation chains. Our core method, chain-of-thought prompting Wei et al. (2022), iteratively breaks down the deduction into simplified logical steps.
@@
Experiments reveal that while zero-shot performance lags due to a lack of grounding, multi-step prompts can unlock substantial reasoning ability in models. Performance improves with more steps and more specificity in the prompt. While introducing few-shot prompting generates good results, we observed that models do significantly better with chain-of-thought prompting. We experiment with several chains of thought and achieve mostly similar results with each attempt. To make a more empirically confident guess towards solving the puzzles we adopt an ensemble of these chains based on majority voting. Our approach achieves competitive performance, ranking 2nd on the word puzzle subtask and 13th on sentence puzzles.
2 Related Work
@@
LLMs have been widely used for complex and challenging language processing tasks recently Raihan et al. (2023a, b); Goswami et al. (2023). They have shown good reasoning abilities in several tasks. The task of solving puzzles and the BrainTeaser dataset Jiang et al. (2023) represent both a novel task and a novel dataset respectively. Similarly to their multiple choice questions (MCQs) approach, a few datasets like MathQA Austin et al. (2021), have been compiled. However, these are intended for specific tasks in which domain knowledge is usually enough thus they not requiring deep reasoning. A similar work is done by Saeedi et al. (2020) where they investigate a task that combines natural language understanding and commonsense reasoning. They present deep learning architectures for distinguishing between sensible and nonsensical statements.
@@
In the field of reasoning task, a chain-of-thought Wei et al. (2022) implies a logical sequence of connected ideas, fostering coherence and depth in responses. On the other hand, a tree-of-thought suggests branching out into various related ideas, offering a more comprehensive exploration of a topic. While few-shot prompting is effective for some tasks by providing examples to guide the model, it may have limitations in capturing the complexity of nuanced conversations. The optimal choice may involve a hybrid approach, where a few-shot prompt sets the initial Yao et al. (2023) context, and the model subsequently follows a chain or tree of thought to generate more contextually rich and coherent responses useful for reasoning tasks.
Tan (2023) shows the performance of LLM’s on the reasoning of arithmetic word problems. It states that higher degrees of realization are associated with better overall accuracy on arithmetic problems. And chain-of-thought is really helpful in this aspect as it covers a variety of prompts to strengthen the reasoning. Similarly, Mo and Xin (2023) presents a new reasoning framework for large language models by addressing a gap in prior tree-based reasoning methods which overlooked inherent uncertainties in intermediate decision points made by models. Overall, the key innovation is leveraging uncertainty estimation locally within the models during tree reasoning to enable more precise problem-solving and reasoning.
3 The BrainTeaser Dataset
The BrainTeaser dataset Jiang et al. (2023), introduced with the task Jiang et al. (2024b) is a question-answering benchmark designed to evaluate models’ ability for lateral thinking, i.e., to defy default commonsense associations and reason creatively. The dataset contains 1,100 multiple-choice questions divided into two sub-tasks - 627 sentence-based puzzles relying on narrative context and common phrases and 492 puzzles focused on the literal form and letters of words.
For a fair comparison with human performance, the dataset also provides a separate human evaluation set with 102 randomly sampled questions. Each question in BrainTeaser has one correct answer and three distractor choices, including the option "none of the above". To prevent memorization of training data, the dataset also contains semantically and contextually reconstructed variants for every question while preserving the original reasoning process and answers. The key statistics of the dataset are shown in Table 1.
Sentence Word
Number of puzzles 627 492
Avg. tokens (prompt) 34.88 10.65
Avg. tokens (choices) 9.11 3.0
Table 1: Key statistics of the BrainTeaser dataset in the sentence and word puzzle sub-task.
During the SemEval-2024 Task 9 development phase, a total of 240 prompts (120 for both sentence and word puzzles) are provided. During the test phase, a total of 216 prompts (120 for sentence and 96 for word puzzles) are provided.
@@ 4 Experiments
In our experiments, we focus on several prompting strategies by employing three state-of-the-art models including proprietary models like GPT-4 OpenAI (2023) and Claude 2.1 Anthropic (2023) (accessed via API key) and one open-source model - Mixtral Jiang et al. (2024a).
@@ 4.3 Chain-of-Thought
4.4 Ensemble of Chain-of-Thought Prompts
@@
This ensemble methodology with chain-of-thought prompting helps improve robustness to outlier examples and noise compared to using a single model. By prompting the ensemble components on different random question subsets, diversity is promoted to capture a more holistic representation of the overall data distribution. The voting also helps cancel out issues with single models latching onto spurious patterns. Overall, the ensemble approach with multiple chain-of-thought prompt sets provides a robust assessment strategy suited for the open-ended nature and diversity of the BrainTeaser puzzles.
@@ 5 Results
We analyze the performance of the three models - including GPT4 Turbo, Claude 2.1, and Mixtral. These models are tested with different types of prompts - regular and chain-of-thought, and with a varying number of examples, known as shots, ranging from zero to eight. Additionally, an ensemble method is applied to the eight-shot chain-of-thought prompting to see if it can further improve the models’ performance. The results, shown in Table 2, reveal how the models performed under each condition. A human baseline with scores of 0.91 for both Sentence and Word puzzles in the test set is provided by the task organizers for comparison purposes.
@@
Table 2: Comparing the results generated by the models with different prompting strategies. [CoT] - denotes chain-of-thought. [E] - denotes Ensemble (as described in 4.4). [*] - denotes submission during the test phase on the Leaderboard.
@@
GPT4 Turbo shows the best performance, especially with chain-of-thought prompting and an increasing number of shots. The model performs best with the eight-shot chain-of-thought prompting combined with the ensemble method ([E]), reaching the highest Sentence and Word scores of 0.93 and 0.95 in the test set, respectively. This shows that chain-of-thought prompting and the ensemble method significantly improve the model’s understanding and output. Claude 2.1 also improves with chain-of-thought prompting and more shots. Its best scores were with the eight-shot chain-of-thought with the ensemble, achieving Sentence and Word scores of 0.86 and 0.95 in the test set, respectively. The asterisk (*) mark in Table 2 denotes our submission during the test phase. Even though Mixtral’s performance is inferior to the performance of the other two models, it consistently gets better with more shots and chain-of-thought prompting. Mixtral delivered best results with the eight-shot chain-of-thought and the ensemble technique, with Sentence and Word scores of 0.88 and 0.82 in the test set, respectively.
@@
Finally, the results highlight the effectiveness of chain-of-thought prompting in boosting the performance of LLMs. This approach, especially when combined with more examples and the ensemble method, greatly improves models’ abilities to process and generate more accurate responses. GPT4 Turbo’s top performance is likely due to its advanced design, which makes the most of these strategies. On the other hand, Claude 2.1’s results point to the importance of model-specific adjustments.
6 Conclusion and Future Work
@@
In this paper, we presented MasonTigers’ approach to SemEval-2024 Task 9 on solving puzzles using LLMs. We explored various prompting strategies to guide the models, including zero-shot, few-shot, and chain-of-thought prompting. Our key method involved iteratively breaking down reasoning into simplified logical steps to decompose the complex deduction process.
@@
Our experiments revealed promising results. While zero-shot performance was limited, providing explanatory prompts substantially improved the models’ reasoning abilities. Performance increased with more prompt specificity and steps. Our best results came from an ensemble approach applying majority voting across multiple chain-of-thought prompts.
@@
Ultimately, our system achieved competitive rankings on the leaderboard, placing $2^{nd}$ in the word puzzle sub-task and $13^{th}$ on sentence puzzles. The strong capability unlocked through guided prompting highlights these models’ latent reasoning potential when given a structured thought process. Our work sheds light on how explanatory chains can elicit more of the knowledge within large language model parameters.
@@
A few key limitations remain to be addressed in future work. First, constructing effective prompts requires extensive human effort and insight - automating this prompting process could improve scalability. Additionally, performance still lags behind human levels, indicating that there is room for advancement. Architectural constraints related to long-term memory and reasoning likely need to be overcome. Finally, our approach focused narrowly on the given puzzles rather than teaching broader inferential skills - developing more generalizable reasoning through prompts is an open challenge.
Acknowledgments
We would like to thank the shared task organizers for proposing this interesting competition and for providing participants with the BrainTeaser dataset.